OpenStudy (0698097):

Question needed to be answered now!! 4x-y=3 7x-2y=11 I need to find the ordered pair of x and y using substitution ive done the problem ten times now im missing something

OpenStudy (mathmale):

Substitution, here, invovles choosing one of the 2 equations and solving it for either x or y. The result is subbed into the other equation. Focus on the first equation. Solve it for y, please. Your answer must look like y=mx+b.

OpenStudy (0698097):

-y=3-4x

OpenStudy (0698097):

y=-3+4x

OpenStudy (mathmale):

Great. Now substitute your (-3+4x) for y in the 2nd equation.

OpenStudy (loveneenah):

It is 3+y/4